# A Stabilized Time-Domain Combined Field Integral Equation Using the Quasi-Helmholtz Projectors

### v1

@@ Abstract
This paper introduces a time-domain combined field integral equation for electromagnetic scattering by a perfect electric conductor. The new equation is obtained by leveraging the quasi-Helmholtz projectors, which separate both the unknown and the source fields into solenoidal and irrotational components. These two components are then appropriately rescaled to cure the solution from a loss of accuracy occurring when the time step is large. Yukawa-type integral operators of a purely imaginary wave number are also used as a Calderón preconditioner to eliminate the ill-conditioning of matrix systems. The stabilized time-domain electric and magnetic field integral equations are linearly combined in a Calderón-like fashion, then temporally discretized using a proper pair of trial functions, resulting in a marching-on-in-time linear system. The novel formulation is immune to spurious resonances, dense discretization breakdown, large-time step breakdown and dc instabilities stemming from non-trivial kernels. Numerical results for both simply-connected and multiply-connected scatterers corroborate the theoretical analysis.
@@ I Introduction
Most individual numerical issues and their combinations have been intensively studied and effectively resolved. For instance, the ill-conditioning of the TD-EFIE can be remedied using a Calderón preconditioning strategy [2, 1], and dc instabilities can be tackled by applying the dot-trick formulation [1] or the loop-tree decomposition [11, 12]. The inaccuracy of the TD-MFIE’s solution has been mitigated by leveraging a mixed discretization scheme [13]. Resonant instabilities are usually eliminated by considering a time-domain combined field integral equation (TD-CFIE) [14, 2, 15]. The low-frequency breakdown and the loss of accuracy can be overcome by means of the loop-star/ loop-tree decompositions, which are particularly effective with simply-connected scatterers [12, 16, 17]. Recently, the quasi-Helmholtz projectors have been introduced in [18] as an alternative to those decompositions and have shown their superiority when applied to both simply-connected and multiply-connected geometries [19, 20].
Nonetheless, a formulation resolving all above mentioned issues is still missing from literature. This deficiency can be explained by the fact that PEC cavities support only discrete interior resonances (i.e., the lowest resonant frequency is far from the very-low-frequency band). Therefore, it does not make sense to consider extremely low frequencies and resonant frequencies in the same regime. However, the loss of accuracy occurs even at moderately low frequencies [21]. In this paper, we propose a stabilized TD-CFIE formulation for moderate to extremely low frequencies, which is immune to resonant instabilities, dense discretization breakdown and large-time step breakdown. For simply-connected geometries, the method is also free from dc instabilities. For multiply-connected geometries, near-dc instabilities are present, but correspond to exponentially decaying error components. Numerical experiments demonstrate they are not problematic in typical applications. The proposed TD-CFIE is a generalization of the quasi-Helmholtz projected (qHP) frequency-domain CFIE introduced in [21].
@@ II-A Time-domain integral equations
Let $\Gamma$ be the surface of a PEC immersed in a homogeneous background medium with permittivity $\epsilon$ and permeability $\mu$. We assume that the surface $\Gamma$ is closed, connected and orientable. The outward unit normal vector to $\Gamma$ is denoted by n. An incident plane wave $(\boldsymbol{e}^{in},\boldsymbol{h}^{in})$ induces a surface current density $\boldsymbol{j}$ on $\Gamma$, which satisfies the following standard TD-EFIE and TD-MFIE:
@@
for all $\boldsymbol{r}\in\Gamma$ and $t>0$. Here, $I$ stands for the identity operator and the time-domain integral operators $\mathcal{T}$ and $\mathcal{K}$ are defined by
@@
where $\eta=\sqrt{\mu/\epsilon},c=1/\sqrt{\mu\epsilon}$, $R=\left|{\boldsymbol{r}-\boldsymbol{r}^{\prime}}\right|$ and $\tau=t-R/c$. In order to guarantee the uniqueness of a time-domain solution $\boldsymbol{j}$, the causality is imposed, i.e., $\boldsymbol{j}(\boldsymbol{r},t)=0$ for all $t<0$ and $\boldsymbol{r}$ in a neighborhood of $\Gamma$.
@@ II-C Spatial discretization
In order to numerically solve boundary integral equations, the surface $\Gamma$ is partitioned into $N_{C}$ planar triangles with $N_{E}$ edges and $N_{V}$ vertices. On the triangular mesh of $\Gamma$, we equip two boundary element spaces of divergence-conforming functions, namely Rao-Wilton-Glisson (RWG) functions $\boldsymbol{f}_{n}(\boldsymbol{r})$ and Buffa-Christiansen (BC) functions $\boldsymbol{g}_{n}(\boldsymbol{r})$, with $n=1,2,\ldots,N_{E}$. The RWG function $\boldsymbol{f}_{n}(\boldsymbol{r})$ associated with the edge $e_{n}$ is defined on two adjacent triangles $c^{+}_{n}$ and $c^{-}_{n}$, which share the common edge $e_{n}$, as follows [24]:
@@
where $A_{c^{+}_{n}}$ and $A_{c^{-}_{n}}$ are the area of $c^{+}_{n}$ and $c^{-}_{n}$, respectively, while $\boldsymbol{r}^{+}_{n}$ and $\boldsymbol{r}^{-}_{n}$ respectively are the vertex of $c^{+}_{n}$ and $c^{-}_{n}$ and opposite to $e_{n}$ (see Fig. 1). In other words, the function $\boldsymbol{f}_{n}(\boldsymbol{r})$ represents a current flowing from $c^{+}_{n}$ to $c^{-}_{n}$. The BC function $\boldsymbol{g}_{n}(\boldsymbol{r})$, which is associated with the edge $e_{n}$, is constructed as a linear combination of the RWG functions defined on the barycentric refinement of the original mesh, with the coefficients specified in [25] (see Fig. 2).
@@
Figure 1: RWG function $\boldsymbol{f}_{n}(\boldsymbol{r})$ associated with the edge $e_{n}$. The function $\boldsymbol{f}_{n}(\boldsymbol{r})$ is defined on two adjacent triangles $c^{-}_{n}$ and $c^{+}_{n}$ sharing the common edge $e_{n}$. The bold arrow indicates the direction of $\boldsymbol{f}_{n}(\boldsymbol{r})$ across $e_{n}$.
@@
Figure 2: BC function $\boldsymbol{g}_{n}(\boldsymbol{r})$ associated with the edge $e_{n}$. The function $\boldsymbol{g}_{n}(\boldsymbol{r})$ is a linear combination of RWG functions defined on the barycentric refinement of the original mesh. The gray area indicates the support of $\boldsymbol{g}_{n}(\boldsymbol{r})$, while the bold arrow indicates the direction of $\boldsymbol{g}_{n}(\boldsymbol{r})$ along $e_{n}$.
The current density $\boldsymbol{j}$ is spatially approximated as an expansion in RWG functions
@@
The TD-EFIE (1) is tested with rotated RWG functions $\textbf{n}\times\boldsymbol{f}_{m}(\boldsymbol{r})$, whereas the TD-MFIE (2) is tested with rotated BC functions $\textbf{n}\times\boldsymbol{g}_{m}(\boldsymbol{r})$ [26, 13], with $m=1,2,\ldots,N_{E}$, resulting in the following semi-discrete equations:
@@
Besides the time-domain integral operators, the Yukawa-type operators are also needed to be spatially discretized. In the next sections, the operators $T_{-j\kappa}$ and $K_{-j\kappa}$ are respectively composed with $\mathcal{T}$ and $\mathcal{K}$. To guarantee the stability of the composition, we discretize $T_{-j\kappa}$ using the BC functions, whereas $K_{-j\kappa}$ is discretized using the RWG trial functions and the rotated BC testing functions [21]. This scheme results in the following discretized matrices:
@@ II-D Quasi-Helmholtz projectors
More specifically, the following quasi-Helmholtz decomposition holds for any RWG coefficient vector $\boldsymbol{i}$:
@@
Unfortunately, the construction of the standard quasi-Helmholtz decomposition is computationally expensive and unstable [30]. The quasi-Helmholtz projectors were introduced in [18] as an alternative to cure the EFIE systems from low-frequency breakdown. More specifically, the quasi-Helmholtz projectors mapping RWG-based functions into irrotational and solenoidal subspaces are respectively defined by
@@
Here, $\mathbb{P}^{\Lambda}$ and $\mathbb{P}^{\Sigma H}$ are the projectors into irrotational and solenoidal spaces, respectively. It is noteworthy that the quasi-Helmholtz projectors do not require an explicit identification of global loops $\operatorname{\mathbf{H}}$. In addition, there is no distinction between local loop and global loop functions.
@@ III-A Plane-wave scattering
When the time step $\Delta t$ is large, the discretized plane-wave incident fields $\mathrm{e}$ and $\mathrm{h}$ scale as follows (see [7, 8, 9, 10, 13]):
@@
Please note that the scaling of $\mathrm{e}$ and $\mathrm{h}$ are not identical because they correspond to different spatial discretization schemes. Next, the induced current density $\mathrm{j}$ has the following large-time step scaling [33]:
@@
where ~$\tilde{\partial}_{t}^{-1}=\partial_{t}^{-1}/T_{0}$ and ~$\tilde{\partial}_{t}=T_{0}\,\partial_{t}$, with $\partial_{t}^{-1}$ the temporal integral over $(-\infty,t)$, $T_{0}=D/c$, and $D$ the diameter of the scatterer. The factor $T_{0}$ introduced in [20] aims to render the “scaled” integral ~$\tilde{\partial}_{t}^{-1}$ and the “scaled” derivative ~$\tilde{\partial}_{t}$ dimensionless (i.e., independent of the chosen units of measurement). Clearly, $\mathrm{y}=\mathcal{O}(1,1,1)^{\mathsf{T}}$, whose components are well-balanced. Irrotational and solenoidal components of the incident electric field $\mathrm{e}$ can also be balanced by using a multiplicative matrix as follows:
@@ III-B Time-domain electric field integral equation
Using the auxiliary unknown $\mathrm{y}$ and the balanced incident electric field $\tilde{\mathrm{e}}$, we obtain the following stabilized TD-EFIE:
@@
Analogously, the discretized matrix $\mathbb{Z}$ of the Yukawa-type EFIE operator $T_{-j\kappa}$, with $\kappa=(c\Delta t)^{-1}$, scales as
@@
Please be aware that the role of $\boldsymbol{\Lambda}$ and $\boldsymbol{\Sigma}$ are exchanged from the RWG basis to the BC basis. Using the quasi-Helmholtz projectors $\mathbb{P}^{\Sigma H}$ and $\mathbb{P}^{\Lambda}$ of the BC functions, we can introduce the following stabilized Yukawa-type EFIE operator:
@@
As $\mathcal{Z}^{h}$, we note that $\mathbb{P}^{\Sigma H}\mathbb{Z}^{h}=\mathbb{Z}^{h}\mathbb{P}^{\Sigma H}=\boldsymbol{0}$. Finally, the matrix $\widetilde{\mathbb{Z}}$ can be used as a Calderón preconditioner of $\widetilde{\mathcal{Z}}$, which results in the following well-conditioned, large-time step well-balanced TD-EFIE:
@@
It is noteworthy that the preconditioner $\widetilde{\mathbb{Z}}$ is independent of time, which saves the cost of additional temporal discretization procedures, as proposed in [20].
@@ III-C Time-domain magnetic field integral equation
In this section, we consider the following symmetrized TD-MFIE, which is inspired by the Calderón-like frequency-domain MFIE formulation in [21, 34]:
@@
The operator $\mathbb{X}\,\mathbb{G}^{-1}\mathcal{X}$ is “symmetric” in the sense that $\mathbb{X}$ and $\mathcal{X}$ are corresponding to the inner and outer MFIE operators, respectively. Since the operators $I$, $\mathcal{K}$ and $K_{-j\kappa}$ are discretized using the same spatial discretization scheme, the semi-discrete operator $\mathcal{X}$ and the matrix $\mathbb{X}$ share the following large-time step scaling (see [8, 10, 13]):
@@
In addition, the algebraic structure of the inverse Gram matrix $\mathbb{G}^{-1}$ is as follows [20]:
@@
Here, the square symbols represent non-zero blocks. Based on those behaviors, the large-time step scaling of the operator $\mathbb{X}\,\mathbb{G}^{-1}\mathcal{X}$, the unknown $\mathrm{j}$ and the right-hand side $\mathbb{X}\,\mathbb{G}^{-1}\mathrm{h}$ in (8) respectively read as:
@@
The symbol $\neq$ indicates that the scaling of the left-hand side and the right-hand side of (8) are not identical. Obviously, at least one of them is incorrect. This discrepancy can be eliminated by taking into account the following property of the static MFIE operators:
@@
where $\mathbb{M}_{0}$ is the discretized matrix of the static operator $K_{0}$. We refer the reader to [21] for a rigorous proof of (9). Besides that, for sufficiently large time step $\Delta t\gg D/c$, the following estimate can be deduced from the Maclaurin series of the exponential function:
@@
Here, the space variables $\boldsymbol{r}$ and $\boldsymbol{r}^{\prime}$ are omitted for the sake of readability. Now, we can rearrange the “loop-loop” components of the operator $\mathbb{X}\,\mathbb{G}^{-1}\mathcal{X}$ as follows:
@@
Consequently, we arrive at the following correct scaling of the operator $\mathbb{X}\,\mathbb{G}^{-1}\mathcal{X}$:
@@
Next, by using appropriate stabilizations for the unknown $\mathrm{j}$ and the right-hand side $\widetilde{\mathbb{X}}\,\mathbb{G}^{-1}\mathrm{h}$, we end up with the following stabilized version of the symmetrized TD-MFIE (8):

### v2

@@ Abstract
This paper introduces a time-domain combined field integral equation for electromagnetic scattering by a perfect electric conductor. The new equation is obtained by leveraging the quasi-Helmholtz projectors, which separate both the unknown and the source fields into solenoidal and irrotational components. These two components are then appropriately rescaled to cure the solution from a loss of accuracy occurring when the time step is large. Yukawa-type integral operators of a purely imaginary wave number are also used as a Calderón preconditioner to eliminate the ill-conditioning of matrix systems. The stabilized time-domain electric and magnetic field integral equations are linearly combined in a Calderón-like fashion, then temporally discretized using an appropriate pair of trial functions, resulting in a marching-on-in-time linear system. The novel formulation is immune to spurious resonances, dense discretization breakdown, large-time step breakdown and dc instabilities stemming from non-trivial kernels. Numerical results for both simply-connected and multiply-connected scatterers corroborate the theoretical analysis.
@@ I Introduction
Most individual numerical issues and their combinations have been intensively studied and effectively resolved. For instance, the ill-conditioning of the TD-EFIE can be remedied using a Calderón preconditioning strategy [2, 1], and dc instabilities can be tackled by applying the dot-trick formulation [1] or the loop-tree decomposition [11, 12]. The inaccuracy of the TD-MFIE’s solution has been mitigated by leveraging a mixed discretization scheme [13]. Resonant instabilities are usually eliminated by considering a time-domain combined field integral equation (TD-CFIE) [14, 2, 15].
The low-frequency breakdown and the loss of accuracy can be overcome by means of the loop-star/loop-tree decompositions, which are particularly effective with simply-connected scatterers [12, 16, 17]. The quasi-Helmholtz projectors have been introduced in [18] as an alternative to those decompositions. Numerous recent efforts leveraging the quasi-Helmholtz projectors to stabilize the low-frequency electromagnetic problems, both in the frequency-domain regime [18, 19, 20, 21] and the time-domain regime [22, 23, 24], have shown their effectiveness when applied to both simply-connected and multiply-connected geometries. In addition to the standard integral equations for the unknown surface electric current, some other approaches have also been proposed to deal with the low-frequency breakdown, such as the current-charge formulation [25], the decoupled potential formulation [26], and the Debye source approach [27, 28].
Nonetheless, a formulation resolving all above mentioned issues is still missing from literature. This deficiency can be explained by the fact that PEC cavities support only discrete interior resonances (i.e., the lowest resonant frequency is far from the very-low-frequency band). Therefore, it does not make sense to consider extremely low frequencies and resonant frequencies in the same regime. However, the loss of accuracy occurs even at moderately low frequencies [19]. In this paper, we propose a stabilized TD-CFIE formulation for moderate to extremely low frequencies, which is immune to resonant instabilities, dense discretization breakdown and large-time step breakdown. For simply-connected geometries, the method is also free from dc instabilities. For multiply-connected geometries, near-dc instabilities are present, but correspond to exponentially decaying error components. Numerical experiments demonstrate they are not problematic in typical applications.
For this purpose, we extend the idea of the quasi-Helmholtz projected (qHP) frequency-domain CFIE introduced in [19]. Firstly, the solenoidal and irrotational components of the unknown and the incident fields are separated using the quasi-Helmholtz projectors and rescaled by appropriate factors, resulting in auxiliary well-balanced unknown and source fields. Secondly, the rescaled TD-EFIE and TD-MFIE are preconditioned by the corresponding stabilized Yukawa-type integral operators of a purely imaginary wave number. Then, their linearly combined formulation is temporally discretized using an appropriate pair of trial functions. The resulting linear system is solved by the MOT algorithm, and a post-processing procedure is performed to obtain the physical unknown.
@@ II-A Time-domain integral equations
Let $\Gamma$ be the surface of a PEC immersed in a homogeneous background medium with permittivity $\epsilon$ and permeability $\mu$. We assume that the surface $\Gamma$ is closed, connected and orientable. The outward unit normal vector to $\Gamma$ is denoted by n. An incident plane wave $(\boldsymbol{e}^{in},\boldsymbol{h}^{in})$ induces a surface current density $\boldsymbol{j}$ on $\Gamma$, which satisfies the following standard TD-EFIE and TD-MFIE
@@
for $\boldsymbol{r}\in\Gamma$ and $t>0$. Here, $I$ stands for the identity operator and the time-domain integral operators $\mathcal{T}$ and $\mathcal{K}$ are defined by
@@
where $\eta=\sqrt{\mu/\epsilon},c=1/\sqrt{\mu\epsilon}$, $R=\left|{\boldsymbol{r}-\boldsymbol{r}^{\prime}}\right|$ and $\tau=t-R/c$. In order to guarantee the uniqueness of a time-domain solution $\boldsymbol{j}$, the causality is imposed, i.e., $\boldsymbol{j}(\boldsymbol{r},t)=\boldsymbol{0}$ for all $t<0$ and $\boldsymbol{r}$ in a neighborhood of $\Gamma$.
@@ II-B Yukawa-type integral operators
Remark II.1.
In the large-time step regime (considered in this paper), the choice $\kappa=(c\Delta t)^{-1}$ (or $\kappa$ of the order of $(c\Delta t)^{-1}$) resolves the large-time step breakdown. In the small-time step regime, other problems occur, and more general choices for the regularizing operators could be appropriate [31, 32].
@@ II-C Spatial discretization
In order to numerically solve boundary integral equations, the surface $\Gamma$ is partitioned into $N_{C}$ planar triangles with $N_{E}$ edges and $N_{V}$ vertices. On the triangular mesh of $\Gamma$, we equip two boundary element spaces of divergence-conforming functions, namely Rao-Wilton-Glisson (RWG) functions $\boldsymbol{f}_{n}(\boldsymbol{r})$ and Buffa-Christiansen (BC) functions $\boldsymbol{g}_{n}(\boldsymbol{r})$, with $n=1,2,\ldots,N_{E}$. The RWG function $\boldsymbol{f}_{n}(\boldsymbol{r})$ associated with the edge $e_{n}$ is defined on two adjacent triangles $c^{+}_{n}$ and $c^{-}_{n}$, which share the common edge $e_{n}$, as follows [33]
@@
where $A_{c^{+}_{n}}$ and $A_{c^{-}_{n}}$ are the area of $c^{+}_{n}$ and $c^{-}_{n}$, respectively, while $\boldsymbol{r}^{+}_{n}$ and $\boldsymbol{r}^{-}_{n}$ respectively are the vertex of $c^{+}_{n}$ and $c^{-}_{n}$ that are opposite to $e_{n}$ (see Fig. 1). In other words, the function $\boldsymbol{f}_{n}(\boldsymbol{r})$ represents a current flowing from $c^{+}_{n}$ to $c^{-}_{n}$. The BC function $\boldsymbol{g}_{n}(\boldsymbol{r})$, which is associated with the edge $e_{n}$, is constructed as a linear combination of the RWG functions defined on the barycentric refinement of the original mesh, with the coefficients specified in [34] (see Fig. 2).
@@
Figure 1: Rao-Wilton-Glisson (RWG) function $\boldsymbol{f}_{n}(\boldsymbol{r})$ associated with the edge $e_{n}$. The function $\boldsymbol{f}_{n}(\boldsymbol{r})$ is defined on two adjacent triangles $c^{-}_{n}$ and $c^{+}_{n}$ sharing the common edge $e_{n}$. The bold arrow indicates the direction of $\boldsymbol{f}_{n}(\boldsymbol{r})$ across $e_{n}$.
@@
Figure 2: Buffa-Christiansen (BC) function $\boldsymbol{g}_{n}(\boldsymbol{r})$ associated with the edge $e_{n}$. The function $\boldsymbol{g}_{n}(\boldsymbol{r})$ is a linear combination of RWG functions defined on the barycentric refinement of the original mesh. The gray area indicates the support of $\boldsymbol{g}_{n}(\boldsymbol{r})$, while the bold arrow indicates the direction of $\boldsymbol{g}_{n}(\boldsymbol{r})$ along $e_{n}$.
The current density $\boldsymbol{j}$ is spatially approximated as an expansion in the RWG basis functions
@@
The TD-EFIE (1) is tested with the rotated RWG functions $\textbf{n}\times\boldsymbol{f}_{m}(\boldsymbol{r})$, whereas the TD-MFIE (2) is tested with the rotated BC functions $\textbf{n}\times\boldsymbol{g}_{m}(\boldsymbol{r})$ [35, 13], with $m=1,2,\ldots,N_{E}$, resulting in the following semi-discrete equations
@@
Besides the time-domain integral operators, the Yukawa-type operators are also needed to be spatially discretized. In the next sections, the operators $T_{-j\kappa}$ and $K_{-j\kappa}$ are respectively composed with $\mathcal{T}$ and $\mathcal{K}$. To guarantee the stability of the composition, we discretize $T_{-j\kappa}$ using the BC functions, whereas $K_{-j\kappa}$ is discretized using the RWG trial functions and the rotated BC testing functions [19]. This scheme results in the following discretization matrices
@@ II-D Quasi-Helmholtz projectors
More specifically, the following quasi-Helmholtz decomposition holds for any RWG coefficient vector $\boldsymbol{i}$
@@
Unfortunately, the construction of the standard quasi-Helmholtz decomposition is computationally expensive and unstable [39]. The quasi-Helmholtz projectors were introduced in [18] as an alternative to cure the EFIE systems from low-frequency breakdown. More specifically, the quasi-Helmholtz projectors mapping RWG-based functions into the irrotational and solenoidal subspaces are respectively defined by
@@
Here, $\mathbb{P}^{\Lambda}$ and $\mathbb{P}^{\Sigma H}$ are the projectors into the irrotational and solenoidal spaces, respectively. It is noteworthy that the quasi-Helmholtz projectors do not require an explicit identification of global loops $\operatorname{\mathbf{H}}$. In addition, there is no distinction between local loop and global loop functions.
@@ III-A Plane-wave scattering
When the time step $\Delta t$ is large, the discretized plane-wave incident fields $\mathrm{e}$ and $\mathrm{h}$ scale as follows (see [7, 8, 9, 10, 13])
@@
Please note that the scaling of $\mathrm{e}$ and $\mathrm{h}$ are not identical because they correspond to different spatial discretization schemes. Next, the induced current density $\mathrm{j}$ has the following large-time step scaling [18]
@@
where ~$\tilde{\partial}_{t}^{-1}=\partial_{t}^{-1}/T_{0}$ and ~$\tilde{\partial}_{t}=T_{0}\,\partial_{t}$, with $\partial_{t}^{-1}$ the temporal integral over $(-\infty,t)$, $T_{0}=D/c$, and $D$ the diameter of the scatterer. The factor $T_{0}$ introduced in [24] aims to render the “scaled” integral ~$\tilde{\partial}_{t}^{-1}$ and the “scaled” derivative ~$\tilde{\partial}_{t}$ dimensionless (i.e., independent of the chosen units of measurement). Clearly, $\mathrm{y}=\mathcal{O}(1,1,1)^{\mathsf{T}}$, whose components are well-balanced. Irrotational and solenoidal components of the incident electric field $\mathrm{e}$ can also be balanced by using a multiplicative matrix as follows
@@ III-B Time-domain electric field integral equation
Using the auxiliary unknown $\mathrm{y}$ and the balanced incident electric field $\tilde{\mathrm{e}}$, we obtain the following stabilized TD-EFIE
@@
Analogously, the discretization matrix $\mathbb{Z}$ of the Yukawa-type EFIE operator $T_{-j\kappa}$, with $\kappa=(c\Delta t)^{-1}$, scales as
@@
Please be aware that the role of $\boldsymbol{\Lambda}$ and $\boldsymbol{\Sigma}$ are exchanged from the RWG basis to the BC basis. Using the quasi-Helmholtz projectors $\mathbb{P}^{\Sigma H}$ and $\mathbb{P}^{\Lambda}$ of the BC functions, we can introduce the following stabilized Yukawa-type EFIE operator
@@
Like $\mathcal{Z}^{h}$, we note that $\mathbb{P}^{\Sigma H}\mathbb{Z}^{h}=\mathbb{Z}^{h}\mathbb{P}^{\Sigma H}=\boldsymbol{0}$. Finally, the matrix $\widetilde{\mathbb{Z}}$ can be used as a Calderón preconditioner of $\widetilde{\mathcal{Z}}$, which results in the following well-conditioned, large-time step well-balanced TD-EFIE
@@
It is noteworthy that the preconditioner $\widetilde{\mathbb{Z}}$ is independent of time, thereby saving the cost of additional temporal discretization procedures, such as the scheme proposed in [24].
@@ III-C Time-domain magnetic field integral equation
In this section, we consider the following symmetrized TD-MFIE, which is inspired by the Calderón-like frequency-domain MFIE formulation in [19, 42]
@@
The operator $\mathbb{X}\,\mathbb{G}^{-1}\mathcal{X}$ is “symmetric” in the sense that $\mathbb{X}$ and $\mathcal{X}$ are corresponding to the inner and outer MFIE operators, respectively. Since the operators $I$, $\mathcal{K}$ and $K_{-j\kappa}$ are discretized using the same spatial discretization scheme, the semi-discrete operator $\mathcal{X}$ and the matrix $\mathbb{X}$ share the following large-time step scaling (see [8, 10, 13])
@@
In addition, the algebraic structure of the inverse Gram matrix $\mathbb{G}^{-1}$ is as follows [24]
@@
Here, the square symbols represent non-zero blocks. Based on those behaviors, the large-time step scaling of the operator $\mathbb{X}\,\mathbb{G}^{-1}\mathcal{X}$, the unknown $\mathrm{j}$ and the right-hand side $\mathbb{X}\,\mathbb{G}^{-1}\mathrm{h}$ in (8) respectively read as
@@
The symbol $\neq$ indicates that the scaling of the left-hand side and the right-hand side of (8) are not identical. Obviously, at least one of them is incorrect. This discrepancy can be eliminated by taking into account the following property of the static MFIE operators
@@
where $\mathbb{M}_{0}$ is the discretization matrix of the static operator $K_{0}$. We refer the reader to [19] for a rigorous proof of (9). Besides that, for sufficiently large time step $\Delta t\gg D/c$, the following estimate can be deduced from the Maclaurin series of the exponential function
@@
Here, the space variables $\boldsymbol{r}$ and $\boldsymbol{r}^{\prime}$ are omitted for the sake of readability. Now, we can rearrange the “loop-loop” components of the operator $\mathbb{X}\,\mathbb{G}^{-1}\mathcal{X}$ as follows
@@
Consequently, we arrive at the following correct scaling of the operator $\mathbb{X}\,\mathbb{G}^{-1}\mathcal{X}$
@@
Next, by using appropriate stabilization for the unknown $\mathrm{j}$ and the right-hand side $\widetilde{\mathbb{X}}\,\mathbb{G}^{-1}\mathrm{h}$, we end up with the following stabilized version of the symmetrized TD-MFIE (8)
@@
Remark III.1.
Even though the low-frequency (large-time step) and dense-mesh regimes share some spectral properties, dense-mesh behaviors of integral operators are more complicated than their large-time step scaling. In this paper, we leverage the Calderón preconditioning technique to eliminate dense-mesh breakdown. This technique is commonly used and thoroughly analyzed in the literature. For the conciseness of the paper, we would then omit the dense-mesh analysis.